OpenStudy (anonymous):

Generate the first 5 terms of this sequence: f(1) = 0 and f(2) = 1, f(n) = f(n - 1) + f(n - 2), for n > 2. 0, -1, 1, 0, 2 0, 1, 1, 2, 3 0, 1, 2, 2, 3 0, 1, 1, 2, 2

OpenStudy (campbell_st):

so n = 3 you need f(3) = f(2) + f(1) = 1 + 0 = 1 n = 4 you need f(4) = f(3) + f(2) = 1 + 1 = 2 n = 5 you need f(5) = f(4) + f(3) = 2 + 1 = 3 and you know f(1) and f(2) so you have the 5 terms